OpenStudy (anonymous):

1. A particle is moving along the curve below. y = 2sin(πx/2) As the particle passes through the point (1/3,1) its x-coordinate increases at a rate of sqrt(10) cm/s. How fast is the distance from the particle to the origin changing at this instant?

OpenStudy (amistre64):

r^2 = x^2+y^2

OpenStudy (amistre64):

x'=sqrt(10) 2r r' = 2x x' + 2y y' r' = (2x x' + 2y y')/2r r' = (x x' + y y')/r we know x'; amd since y is a function of x we can know y' as well

OpenStudy (amistre64):

y = 2sin(πx/2) y' = pi cos(x pi/2)